# Analysis of optical pattern formation on glass

### v1

@@ Abstract
We investigate the formation of particular patterns when light passes through glass. Experimentally, we use various glass plates, registering each reflected and transmitted outcome. Thus, we find the condition to form a common pattern, namely, randomly scratched plates produce the halos. We engineer other optical patterns by means of the specular holography method. Using the Defocusing Microscopy technique, we study the geometric properties of the glass, establishing the physical and mathematical model for a simulation. Our computational results agree well with the experimental ones, leading us to conclude that the pattern formation is governed by specular reflection on various points along the scratches. Our findings give an initial explanation for the pattern formation on glass, in particular for the halo design.
@@ 1 Introduction
In optics, the interaction of light with rough surfaces is fundamental to understanding everyday effects, considering that objects are not perfectly smooth. For instance, the light scattering phenomenon involves the redirection of light by oscillating charges and it approaches the limit of geometric optics for large particles in comparison to the wavelength of light [1]. The scattered light intensity and profile were predicted by various complex methods [2, 3, 4], but results reported in the literature indicate that the theory is not yet complete [5, 6]. Interestingly, the surface roughness can be used to create holograms. Experiments with plastic paper and a compass revealed a technique, called specular holography, for creating holograms by hand [7, 8]. The arched scratches on the surface act as a curved mirror and form the tridimensional (3D) image. A computational method successfully predicted the bright spots that compose the hologram by applying geometric optics and specular reflection [9], which greatly simplify the light-surface interaction in this scope.
@@
Previous works mainly focused on developing broad theories for the interaction between light and surface roughness. However, the unguided formation of certain patterns on the surface of reflective materials, mainly glass, remains to be described. In this paper, we start from the problem “Glass halo”, proposed for the International Physicists’ Tournament (IPT) 2023 [10]. The problem statement goes as: “Glittering circles can be seen when light from a source with small angular size passes through a glass. On closer examination they appear to be composed of small scratches and structural inhomogeneities. In some cases, specific rays can be seen, diverging from the light source. Under which conditions such circle halos and lines can be seen? Investigate their geometrical properties and what shapes you can engineer.” Accordingly, we started by experimenting with microscope glass plates and registering the pattern outcome. A geometric investigation laid the basis for a simulation that agrees with the geometric optics simplification proposed by [9]. As we will show, the resulting optical design allows us to specify the roughness pattern in various materials. This identification is an initial progress when dealing with everyday objects that do not always appear rough to the human eye; for instance, the patterns discussed here can be seen in glass windows, cellphone screens and car surfaces. From the viewpoint of applications, the understanding of pattern formation associated with roughness improves studies of surface polishing.
@@
The remaining sections are organized as follows. In Sec. 2, we present all methods: experimental procedures with glass plates, simulation, and experimental process with the specular holography technique. In Sec. 3, we present and discuss the experimental results, comparing them to the simulation when suitable, and discuss the points for additional studies. Finally, our concluding remarks are stated in section 4.
@@ 2.1 Experiments
We started our investigation with experiments varying two factors: (a) the glass surface and (b) the distance between the glass plate and the camera-light source combination: see Fig. 1a for our experimental setup. We used a polished as well as mechanically scratched glass plates. The materials and their specifications are listed in Table 1. The cellphone flashlight is composed of light emitting diodes (LEDs) and its intensity was constant during all the experiments carried out for this paper.
@@
During the first part, we took photos of the reflected (observer 1) and transmitted (observer 2) patterns for each glass plate (see again Fig. 1). In all these cases, the distance between the glass surface and the flashlight was $d_{f}=(0.55\pm 0.01)$m. Similarly, the distance between the surface and both cameras (observers) was always $d_{c}=(0.55\pm 0.01)$m.
In the second round of experiments, we varied the distances $d_{f}$ and $d_{c}$ for plates 5 and 6, taking photos of the reflected and transmitted patterns. Starting at $d_{f}=d_{c}=(0.55\pm 0.01)$m, we reduced ten centimeters at a time until $d_{f}=d_{c}=(0.25\pm 0.01)$m.
Figure 1: (a) Sketch of the experimental setup. The distances from the glass plate to the flashlight $d_{f}$ and to the camera $d_{c}$ at either positions (viewer 1 or 2) were equal. (b) Glass plates described in Table 1. Since the light source is not perpendicular to the scratches, we do not see the formation of optical patterns.
Subsequently, we studied the geometrical properties of the glass plates in order to characterize the scratches properties. In this context, the defocusing microscopy technique [11] is used to reconstruct the thickness and curvature profiles of phase objects, such as red blood cells [12] and waveguides [13]. We used it to reconstruct the profile of the scratches on selected regions of plates 4, 5, and 6. For cases that only involve refraction and reflection, the use of geometric optics is more than sufficient. Thus, we checked whether the scratches have dimensions greater than the wavelength of light.
@@ 2.2 Simulation
to plot the light rays. In equation (1), $\hat{E}(t)$ is the unit vector in the direction of the incoming ray, that is, corresponding to the vector that goes from the light source to a given point on the scratch, while $\hat{R}(t)$ defines the unit vector in the direction of the outgoing ray, i.e., the direction of the vector that goes from the point on the scratch to the camera (see Fig. 2). The unit vector $\hat{q}(t)$ defines the tangent line to the scratch curve.
Figure 2: Simulation outline. The incoming ray reaches the scratch and is reflected towards the camera. The reflection is seen (plotted) if it obeys the tolerance criterion.
The scratches on everyday objects are random unless it undergoes an ordered process. Also, we assume that the scratches behave like a gap in the glass with semicircular cross section and are locally linear. This shape produces the following feature: the reflected ray can exit at any angle in the plane passing through the section, since it depends on the point that the light hits the semicircle. This implies that a scratch can reflect the ray at any angle in a plane passing through the point, which is a reasonable assumption, considering that we do not know the profile of the scrape. Lastly, the light source is considered point-like.
Equation (1) can be written as
$|\hat{E}(t)||\hat{q}(t)|cos(\theta_{1})=|\hat{R}(t)||\hat{q}(t)|cos(\theta_{2}),$ (2)
where $\theta_{1}$ is the angle between $\hat{E}(t)$ and $\hat{q}(t)$ and, similarly, $\theta_{2}$ is the angle between $\hat{R}(t)$ and $\hat{q}(t)$. For the camera or our eyes to capture the reflected beam the following relation must be satisfied, within a narrow tolerance,
@@
In our simulation, written in Python, first we generate the scratches (randomly or ordered) and then test the law of reflection for every point along the scratches. Only points that obey the cosine difference of Eq. (3) for a given tolerance ($TOL=0.0125$) are visible in the plotted results.
@@ 3.1 Observed patterns
First, the randomly scratched glass plates generate the glass halo, as shown in Fig. 4 for plates 4, 5, and 6 described in Table 1. The halos are visible for all the randomly scratched plates regardless of the sandpaper used, and, as expected, the halo lines appear more fine as we increase the sandpaper grit. We also observed the formation of this halo in other materials with randomly scraped surfaces, such as glass windows and cellphone screens. This type of surface reflects the light from all directions, and a circular design is anticipated.
@@
The second main pattern is the formation of diverging lines, as seen in Fig. 5. They are an optical distortion in the imaging process and it can occur to any of the plates of Table 1, since it is related to the way the camera lenses capture the strong light source. Indeed, in our experiments, we noticed that the lines are not visible to the bare eyes and that they rotate as the camera rotates (but not when the flashlight or the glass plate rotates).
@@ 3.2 Scratch reconstruction
If we consider a scratched plate on the xy plane and look at it from the z axis, the width of the scratch on this plane determine the size of the obstacle/slit. Therefore, the scratch’s width is the relevant parameter to compare with the wavelength of light.
Using the defocusing microscopy method, we estimated the mean width $a$ of the scratches after reconstructing their surface and lateral profile, see Fig. 6. From the results presented in Table 3, we observe that the width tends to be larger than the wavelength range of the cellphone flashlight $\lambda_{\,light\,}=(0.400-0.750)\mu$m [15]. Also, in our experiments, we do not observe interference or diffraction patterns. These factors indicate that the geometric optics hypothesis, assumed in our simulation, is reasonable.
Plate Estimated size
Plate 4 $a=(15\pm 2)\mu$m
Plate 5 $a=(15\pm 2)\mu$m
Plate 6 $a=(16\pm 6)\mu$m
Table 3: Mean width of the reconstructed scratches for the randomly scratched plates. The uncertainties are the standard deviation from seven measurements.
Figure 6: (a) Reconstruction of plate 5. We obtained the mean width from the marked scratches. (b) Three dimensional plot of a selected reconstructed area. (c) Lateral profile of a selected scratch.
@@ 3.3 Simulation
Starting with the horizontally and vertically scratched glass (plates 2 and 3), in Fig. 7 we show a comparison between the experimental and the theoretical results. In the results, we notice that a horizontally scraped plate highlights a vertical light pattern, and vertically scratched plates produce a horizontal highlight. We see a close similarity between the predicted and the experimental patterns.
@@
Proceeding to the case of randomly scratched plates, we first compare the simulated and experimental results in Fig. 8, confirming that our model describes well our results.
The consistency between the results indicate that the simplifications of our model are valid, and we used our simulation to study a characteristic of the phenomenon, i.e., the halo density. We did it by altering two factors: the number of scrapes and the distance to the light-camera combination.
As we progressively increase the number of scratches on the surface, the halo density also increases, as we can observe in Fig. 9. It happens because for a larger number of scratches, more points obey the law of reflection for the small tolerance imposed, and the halo will appear fuller.
Similarly, for a fixed number of scratches, we see a fuller circle of halos when the plate is relatively far from the light-camera group - see now Fig. 10. As we move the surface closer to the light and the camera, fewer light rays contact the scrapes and are reflected in the desired angle of the observer.

### v2

@@ Abstract
Motivated by a problem from the 2023 International Physicists’ Tournament, we investigate the formation of particular patterns when light passes through glass. Experimentally, we use various glass plates, registering each reflected and transmitted outcome. Thus, we find the condition to form a common pattern, namely, randomly scratched plates produce the halos. We engineer other optical patterns by means of the specular holography method. Using the Defocusing Microscopy technique, we study the geometric properties of the glass, establishing the physical and mathematical model for a simulation. Our computational results agree well with the experimental ones, leading us to conclude that the pattern formation is governed by specular reflection on various points along the scratches. Our findings give an initial explanation for the pattern formation on glass, in particular for the halo design. We also show that the IPT open questions help students learn physics concepts and develop research skills.
@@ 1 Introduction
In optics, the interaction of light with rough surfaces is fundamental to understanding everyday effects, considering that objects are not perfectly smooth. For instance, the light scattering phenomenon involves the redirection of light by oscillating charges, and it approaches the limit of geometric optics for large particles in comparison to the wavelength of light [1]. The intensity and profile of the scattered light are studied using complex methods and approximations, as summarized in Refs. [2, 3, 4, 5, 6]. We can divide these methods into two main categories: (i) in analytical theories [7, 8], certain approximations are used to solve Maxwell’s equations, which have limitations and a domain of validity, while (ii) in numerical models [9, 10] Maxwell’s equations are solved in differential or integral form. Various parameters and regimes can be analyzed: small to large roughness, different scattering materials, and roughness distributions. Thus, results are usually reported based on specific cases, both for analytical [11, 12, 13, 14] and numerical [15, 16, 17] studies.
@@
These mentioned works concentrated on developing broad theories for the interaction between light and surface roughness, not focusing on the formation of optical patterns. Interestingly, surface roughness can be used to create holograms. Experiments with plastic paper and a compass revealed a technique, called specular holography, to create holograms by hand [18, 19]. Arched scratches on the surface act as a curved mirror and form a tridimensional (3D) image. A computational method successfully predicted the bright spots that make up the hologram by applying geometric optics and specular reflection [20], which greatly simplifies the light-surface interaction in this scope. Thus, the relation of roughness and pattern formation started to be explored in the holographic technique.
Even so, the unguided formation of certain patterns on the surface of reflective materials, mainly glass, remains to be described. In this paper, we start from the problem “Glass halo”, proposed for the International Physicists’ Tournament (IPT) 2023 [21]. The IPT is the largest physics competition between teams of university students from different countries. Each year, 17 open problems are proposed, and the teams have a couple of months to prepare their solutions. During the tournament rounds, a team presents and debates their solution with other teams [22]. These open problems provide a research prospect to undergraduate students, allowing them to creatively elaborate experiments and apply physics concepts while working as a team. Thus, when working on IPT problems, students develop relevant abilities to their academic career, such as research planning and effective communication. In particular, we worked on this problem while preparing for the Brazilian Physicists’ Tournament (BPT), which follows the same format and selects the Brazilian team that will participate in the next IPT. During this research process, we explored various possibilities and debated with our colleagues and professors until reaching the solution detailed in this paper. Our results were presented in the 2022 BPT and the 2023 IPT Conference; for the first time the conference abstracts were published as proceedings [23].
@@
The “Glass halo” problem statement goes as: “Glittering circles can be seen when light from a source with small angular size passes through a glass. On closer examination they appear to be composed of small scratches and structural inhomogeneities. In some cases, specific rays can be seen, diverging from the light source. Under which conditions such circle halos and lines can be seen? Investigate their geometrical properties and what shapes you can engineer.” Accordingly, we started by experimenting with microscope glass plates and registering the pattern outcome. A geometric investigation laid the basis for a simulation that agrees with the geometric optics simplification proposed by [20]. As we will show, the resulting optical design allows us to specify the roughness pattern in various materials. This identification is an initial progress when dealing with everyday objects that do not always appear rough to the human eye; for instance, the patterns discussed here can appear in glass windows, cellphone screens and many other reflective surfaces, as seen in Fig. 1. From the viewpoint of applications, the understanding of pattern formation associated with roughness improves studies of surface polishing.
Figure 1: Halo pattern seen on (a) a spoon, (b) a car surface and (c) a ruler.
Our goal is (i) to present the steps taken to arrive at our solution for this IPT problem and (ii) describe the benefit of these open questions in teaching physics concepts and scientific methods. For example, while tackling the Glass Halo problem, we planned experiments and pursued an inquiry-based learning, since we could not find an explanation for it in the literature and had to test various scenarios to find the right conditions. Additionally, our solution can now help undergraduate students visualize optical concepts with an accessible experiment and learn geometric optics simulation.
@@
The remaining sections are organized as follows. In Sec. 2, we present all methods: experimental procedures with glass plates, simulation, and experimental process with the specular holography technique. In Sec. 3, we present the experimental results, comparing them to the simulation when suitable; we also discuss the points for additional studies along with the pedagogical application of our results. Finally, our concluding remarks are stated in section 4.
@@ 2.1 Experiments
We started our investigation with experiments varying one parameter at a time: (a) first the glass surface and (b) then the distance between the glass plate and the camera-light source combination: see Fig. 2a for our experimental setup. We used polished as well as mechanically scratched glass plates. The materials and their specifications are listed in Table 1. The cellphone flashlight is composed of light emitting diodes (LEDs) and its intensity was constant during all the experiments carried out for this paper. Also, the glass surface is always parallel to the cellphones.
@@
During the first part, we took photos of the reflected (viewer 1) and transmitted (viewer 2) patterns for each glass plate (see again Fig. 2). In all these cases, the distance between the glass surface and the flashlight was always $d_{f}=(0.55\pm 0.01)\,$m. Similarly, the distance between the surface and both cameras (observers) was always $d_{c}=(0.55\pm 0.01)\,$m to avoid obstructions and shadows in the way.
In the second round of experiments, we varied the distances $d_{f}$ and $d_{c}$ for plates 5 and 6, taking photos of the reflected and transmitted patterns each time. Starting at $d_{f}=d_{c}=(0.55\pm 0.01)\,$m, we reduced ten centimeters at a time until $d_{f}=d_{c}=(0.25\pm 0.01)\,$m.
Figure 2: (a) Experimental setup. Inset: Sketch showing the important parameters. For illustrative purposes only. (b) Glass plates described in Table 1. Since light is not hitting the glass surface perpendicularly (it is coming from above), we do not see the formation of optical patterns here.
2.2 Scratch reconstruction
Subsequently, we studied the geometrical properties of the glass plates in order to characterize the scratches properties. In this context, the Defocusing Microscopy (MD) technique [24] is used to reconstruct the thickness and curvature profiles of phase objects, such as red blood cells [25] and waveguides [26]. The basic idea behind this technique is that phase objects can be observed in bright-field microscopes when defocused.
That means this method could be used to reconstruct the profile of the scratches on selected regions of plates 4, 5, and 6, since they are phase objects, which otherwise would be transparent under bright-field microscopes. In MD, the image contrast for the scratch is given by
$C(z_{f1})=-\frac{I-I_{0}}{I_{0}}=\frac{\Delta n}{n_{o}}\,z_{f1}\,\nabla^{2}h,$ (1)
II being the intensity of the image, $I_{0}$ the mean intensity of the image, $\Delta n$ being the difference between the refractive index of air and the plate glass, $n_{o}$ the refractive index of the objective immersion fluid (in our case, $n_{0}=1$ since we use a air objective), $z_{f1}$ the distance to a reference plane and $h$ the object profile. The equation show us that the contrast of these images is proportional to the Laplacian of the objects surface, in other words, to the local curvature of the sample.
The technique consists in selecting an area of interest of the plate and adjusting the microscope to obtain an image with the least possible contrast as a reference plane $z_{f0}$. After that, two images are captured, one defocused above $(z_{f1})$ and the other below $(z_{f2})$. These images, each having different contrasts $C(z_{f1})$ and $C(z_{f2})$, can be subtracted from one another, pixel by pixel, to create a third image representing the difference of contrast between the two, described by:
$C(z_{f2})-C(z_{f1})=\Delta n\,(z_{f2}-z_{f1})\,\nabla^{2}h,$ (2)
Since we want the scratch profile $h$, we have:
$\nabla^{2}h=\frac{C(z_{f2})-C(z_{f1})}{\Delta n\,(z_{f2}-z_{f1})},$ (3)
This is a Poisson equation that can be solved in the Fourier space and the solution is given by
$\Delta h(x,y)=\frac{p^{2}}{\Delta n(z_{f2}-z_{f1})N^{2}}\mathcal{F}^{-1}\left\{\frac{\mathcal{F}\{C(z_{f2})-C(z_{f1})\}}{q\,^{2}(x)+q\,^{2}(y)}\right\},$ (4)
where $\mathcal{F}$ is the Fourier transform, $q(x)$ and $q(y)$ are the spatial frequencies of the image, $N$ the size of the image window and $p$ the size of each pixel in micrometers.
The scratched plates were analyzed in an inverted microscope Nikon Eclipse TiE using $10X$ (NA 0.55) and $40X$ (NA 0.25) objectives and a Silicon Video SV643M CMOS camera. We can estimate the lateral resolution of our images. For the 40x(NA 0.55) objective, the resolution is around $0.75\,\mu m$ and, for the 10x (NA 0.25) one, the resolution is around $1.10\,\mu m$. As for the measurements of the scratches depth in the direction of the light propagation, we have an increased resolution, since defocusing microscopy measures the phase difference of the light and is an interferometric technique, similar to other optical phase microscopy techniques. To estimate the sensitivity to measure the depth of the scratches, we can follow equation 10 in [25], which gives for a scratch with $10\,\mu m$ radius a 100nm100\,nm sensitivity.
@@ 2.3 Simulation
to plot the light rays. In equation 5, $\hat{E}(t)$ is the unit vector in the direction of the incoming ray, that is, corresponding to the vector that goes from the light source to a given point on the scratch, while $\hat{R}(t)$ defines the unit vector in the direction of the outgoing ray, i.e., the direction of the vector that goes from the point on the scratch to the camera (see Fig. 3). The unit vector $\hat{q}(t)$ defines the tangent line to the scratch curve. For a generic scratch curve, equation 5 is not able to be solved analytically, and therefore needs root finding algorithms that naturally introduce a narrow tolerance for the equation. Numerically, what the computer does is finding values of $t$ that satisfy:
$\left|\hat{E}(t)\cdot\hat{q}(t)-\hat{R}(t)\cdot\hat{q}(t)\right|\leq TOL,$ (6)
and since $|\hat{E}(t)|=1$, $|\hat{R}(t)|=1$ and $|\hat{q}(t)|=1$, the camera or our eyes will capture the reflected beam if the following relation is satisfied
@@
Figure 3: Simulation outline. The plane containing the light source and the detector is parallel to the plate, as seen in Fig. 2; here the image is rotated to visualize better. The incoming ray reaches the scratch and is reflected towards the camera. The reflection is seen (plotted) if it obeys the tolerance criterion.
Also, scratches on everyday objects are random unless they undergo an ordered process. We assume that the scratches behave like a gap in the glass with semicircular cross section and are locally linear. This shape produces the following feature: the reflected ray can exit at any angle in the plane passing through the section, since it depends on the point that the light hits the semicircle. This implies that a scratch can reflect the ray at any angle in a plane passing through the point, which is a reasonable assumption, considering that we do not know the profile of the scrape. Lastly, the light source is considered point-like. See A.1 for a detailed explanation of these assumptions and an example.
In our simulation, written in Python and available at [27], first we generate the scratches (random or ordered) and then test the law of reflection for every point along the scratches. Only points that obey the cosine difference of Eq. 7 for a given tolerance ($TOL=0.0125$) are visible in the plotted results.
@@ 2.4 Specular Holography
Thus, we see that the IPT problems encourage students to experiment and find the relevant parameters first; later, students can strengthen their knowledge through more advanced techniques and topics. Initially, we performed various experiments to observe the patterns and propose the methods described in Section 2.1. Then, we were able to research different approaches to explain it, arriving at the Defocusing Microscopy technique. Additionally, we could expand our knowledge on optics simulations using Python and on holograms.
@@ 3.1 Observed patterns
First, the randomly scratched glass plates generate the glass halo, as shown in Fig. 5 for plates 4, 5, and 6 described in Table 1. The halos are visible for all the randomly scratched plates regardless of the sandpaper used. As expected, the halo lines appear finer as we increase the sandpaper grit, while a coarse sandpaper produces a rougher and undefined circular pattern. We also observed the formation of this halo in other materials with randomly scraped surfaces, such as glass windows, cellphone screens, cars, spoons, and reflective surfaces in general (Fig. 1). This type of surface reflects light from all directions and a circular design is anticipated.
@@
The second main pattern is the formation of diverging lines, as seen in Fig. 6. They are an optical distortion in the imaging process and it can occur to any of the plates of Table 1, since it is related to the imaging process [29]. Indeed, in our experiments, we noticed that the lines are not visible to the bare eyes and that they rotate as the camera rotates (but not when the flashlight or the glass plate rotates). Researching the available literature along with experimentally checking the known results are a crucial part of tackling an IPT problem. In particular, this distortion is well-established in photography [29] and we have probed its behavior in our experimental setup, e.g. by rotating the experimental components, to check if we were really dealing with a distortion caused in the imaging process.
@@ 3.2 Scratch reconstruction
The halo pattern is just seen on the plates that were scratched with sandpaper, implying that the structures that generate this phenomenon are created in the process of scratching. At first, it is reasonable to assume that the scratch width could have any size, ranging from much smaller to much bigger than the visible light wavelength. However, the halo effect and scratches can be seen with the naked eye, which indicates that scratches with widths bigger than the visible light wavelength do exist. Most importantly, it can be seen during the experiments that parts of these visible scratches “shine” to help create the halo effect, as seen in Fig. 5 for example, which indicates that these larger scratches must be taken into account to describe this phenomenon. It does not mean that other smaller structures do not exist: the ones with similar size, but larger than the wavelength would create diffraction or interference patterns, but these patterns or fringes are not observed in our microscopy experiment, while smaller structures would generate evanescent waves that would not propagate to an observer. Therefore, our object of interest are the structures with a width greater than the visible light wavelength.
Using the Defocusing Microscopy technique, we were able to reconstruct the topology of parts of the randomly scratched plates (4, 5 and 6) and measure the mean width and depth of the scratches on a plate, see Fig. 7 (a). We also calculated the standard deviation of each quantity to measure how much the size distribution varies in each plate, see now Fig. 7 (c).
Taking in account the size of the structures and that they do not present a diffraction or interference pattern we assumed a geometric optics approach in the simulation to replicate the phenomena.
Plate (Nº of samples) Mean width Width SD
[…]
Table 3: Mean width and depth of the reconstructed scratches. Data available at [27].
Figure 7: (a) Example of reconstruction of one of the plates (plate 5). The marked lines were used to calculate the mean width and depth. (b) Width distribution of all the 45 scratch samples collected in plates 4, 5 and 6. (c) Example of a lateral profile of a selected scratch.
@@ 3.3 Simulation
Starting with the horizontally and vertically scratched glass (plates 2 and 3), in Fig. 8 we show a comparison between experimental and theoretical results. In the results, we notice that a horizontally scraped plate highlights a vertical light pattern, and vertically scratched plates produce a horizontal highlight. We see a close similarity between the predicted and the experimental patterns.
@@
Proceeding to the case of randomly scratched plates, we first compare the simulated and experimental results in Fig. 9, confirming that our model describes our results well. The consistency between the results indicate that the simplifications of our model are valid, and we can now use it to study other aspects of the phenomenon. The simulation becomes particularly useful when we lack the precision to perform a quantitative analysis - see Sec. 3.5. Thus, we studied the halo density by altering two factors in the simulation: the number of scrapes and the distance to the light-camera combination.
As we progressively increase the number of scratches on the surface, the halo density also increases, as we can observe in Fig. 10. This happens because for a larger number of scratches, more points obey the law of reflection for the small tolerance imposed, and the halo will appear fuller.
Similarly, for a fixed number of scratches, we see a fuller circle of halos when the plate is relatively far from the light-camera group - see now Fig. 11. As we move the surface closer to the light and the camera, fewer rays contact the scrapes and are reflected within the desired angle of the observer. These results are intuitive and confirm than our model can reproduce the phenomenon.